2) What volume of 50 mM Na2HPO4 and 100 mM NaH2PO4 is needed to prepare 3L of 25 mM
phosphate buffer, pH 6.5 (pKa = 7.21)?
1
Expert's answer
2021-08-30T17:16:02-0400

Using Henderson equition

"PH=pKa+log\frac{A^-}{HA}\\6.5=7.21+log\frac{A^-}{HA}\\log\frac{A^-}{HA}=0.1950\\^-=0.1950(HA)"

Also

"(A^-)+(HA)=25mm\\0.1950(HA+(HA)=25mm\\(HA)=20.9208mm\\(A^-)=0.1950(HA)\\=4.0792"

Need: 3 L of 20.9206mm NaH2PO4


"M_1V_1=M_2V_2\\100mm×V=20.9206mm×3L\\V=0.6276L\\=627.624ml"

Need: 3 L of 4.0792mm NaH2PO4


Similarly

"50mm×V=4.0792mm×300ml\\=244.75mL"